# Connecting the Dots: A Chain-of-Collaboration Prompting Framework for LLM Agents

### v1

@@ 4.2 Enhancing Collaborative Reasoning Capabilities in LLM Agents with Cochain
Table 4: Results of multi-stage collaboration performance on Automotive datasets, highlighting fine-tuned results, Cochain application, and improvements()($\uparrow$)( ↑ ). The experimental results reflect two different scenarios in the workflow. The score is derived from GPT-4’s rating of the final solution (10-point scale), which is synthesized by integrating the answers from each stage.
Stage Method Backbone BS-F GLEU ROUGE-L Backbone BS-F GLEU ROUGE-L
$S_{1}$ Finetuned Qwen2-7B 71.60 22.59 24.51 Llama2-7B 70.68 18.79 21.49
+ Cochain 74.65 26.28 25.97 71.46 20.96 21.58
$\uparrow$ +3.05 +3.69 +1.46 +0.78 +2.17 +0.09
$S_{2}$ Finetuned 70.44 19.37 21.70 GLM4-9B 69.34 16.30 19.76
+ Cochain 74.65 26.57 25.74 73.19 22.33 24.06
$\uparrow$ +4.21 +7.20 +4.04 +3.85 +6.03 +4.30
$S_{3}$ Finetuned 70.23 19.39 22.39 Qwen2-7B 70.28 18.95 22.38
+ Cochain 74.84 27.39 26.29 74.51 25.76 26.00
$\uparrow$ +4.61 +8.00 +3.90 +4.23 +6.81 +3.62
$S_{4}$ Finetuned 69.42 18.71 21.90 DeepSeek-7B 68.96 13.06 18.70
+ Cochain 75.39 29.28 27.94 71.03 14.39 18.89
$\uparrow$ +5.97 +10.57 +6.04 +2.07 +1.33 +0.19
Score 6.3 9.2 +2.9 6.5 9.6 +3.1
4.3 Cochain Efficacy Across Specialized and Cross-Domain Tasks
To investigate Cochain’s impact on specialized and cross-domain tasks, we conducted experiments on automotive industry data using two approaches: a multi-domain model (single Qwen2-7B fine-tuned on all domains) and specialized models (different models each fine-tuned for specific domains), with testing conducted on 4 sub-datasets. Table 4 shows that Cochain significantly improves both approaches, with particularly strong GLEU score gains. The specialized configuration demonstrates Cochain’s ability to orchestrate effective collaboration among domain-expert agents. GPT-4 evaluation scores, assessing comprehensive solutions resulting from four stages, confirm substantial quality improvements for both approaches. Additional experiments with specialized models of the same architecture fine-tuned for different domains are presented in Appendix C.3, further validating Cochain’s effectiveness in coordinating expert agents toward coherent solutions.
Table 5: Result of ablation study. $\mathcal{KG}$ represents the knowledge graph. We highlight the most and second-most efficient modules.
@@ 4.3 Cochain Efficacy Across Specialized and Cross-Domain Tasks
4.4 Key Design Factors and Hyperparameter Influence in Cochain
@@
To validate Cochain’s effectiveness, we conduct ablation experiments with three variants: (1) w/o Knowledge Graph; (2) w/o Causal Chain; and (3) w/o Prompts Tree. As shown in Table 5, all variants exhibit performance degradation across all metrics, confirming each component’s importance. Notably, removing the prompts tree causes the most significant drop, highlighting its critical role in collaboration(Appendix C.4). The knowledge graph and causal chain components also prove essential, demonstrating their complementary functions in the reasoning process.
Hyper-parameter Sensitivity Analysis.
We investigated key hyperparameters by fixing the number of nodes ($n$) from the knowledge graph while varying the prompt count. Figure 7 shows that for any fixed $n$, performance follows an inverted U-shaped curve, indicating an optimal range rather than a monotonic relationship. When fixing prompt count and varying $n$, no consistent pattern emerged, suggesting prompt design exerts greater influence than knowledge quantity, aligning with our ablation results. Table 3 demonstrates that skipping any stage ($S_{k}$) in the prompts tree consistently reduces performance, confirming each stage’s indispensable role in the collaborative reasoning process.
4.5 Overcollaboration is More Serious than Undercollaboration
Business workflow tasks are particularly susceptible to the “overcollaboration” phenomenon. When multiple agents collaborate, this manifests as responses deviating from core issues and reducing answer quality. Comparative analysis of Table 1, Table 4, and additional single-agent experimental results (Appendix C.2) demonstrates that, despite consuming significantly more computational resources, overcollaboration performs substantially worse than undercollaboration. Overcollaboration’s challenging nature is its covertness—systems maintain high activity and apparent collaboration, making the problem hard to detect and rectify promptly. While task decomposition, such as approaches like PMC, and summarization, such as strategies like CoA, are considered effective methods for focusing on core issues [61, 14], Cochain’s collaboration mechanism exhibits superior performance in controlling excessive collaboration, enabling a more precise focus on critical task requirements.
@@ 5 Conclusion
In this paper, we propose Cochain, a chain-of-collaboration prompting framework designed to enhance the collaboration of LLM agents in workflow tasks. Cochain combines the advantages of both single-agent and multi-agent systems, achieving depth in thought and breadth in collaboration. It is characterized by low operational cost, high inference speed, strong interpretability, and robust reasoning chain resilience. Cochain introduces a prompts tree structure for reasoning, enabling collaborative inference paths among multiple agents on-chain. By constructing a cross-node tacit knowledge graph, it improves relational reasoning and reduces hallucinations during collaboration. Additionally, the causal chain mechanism enhances interpretability. Extensive experiments demonstrate that Cochain yields effective improvements in scenarios suffering from both undercollaboration and overcollaboration, excelling in specialized and cross-domain tasks. Expert evaluations also indicate that combining smaller models with Cochain surpasses GPT-4 in workflow tasks.
Appendix A Proof: Tacit Knowledge & Counterfactual Output Model
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Appendix B Implementation Details
Table 6 provides a detailed overview of the datasets, which include three business workflow datasets comprising eleven subsets. Notably, the automotive datasets used for fine-tuning and evaluation are meticulously calibrated by experts from a major automotive manufacturing enterprise.
@@ Appendix B Implementation Details
Table 6: Dataset Statistics. The table shows detailed information about our three domain-specific business workflow datasets. Seed sources indicate the origin of initial data from which the complete datasets were generated using the Self-Instruct [55] methodology.
Datasets Automotive Pharmaceutical E-commerce
Design Production Supply QC R&D Mfg. Sales Procure. Logistics Operations UX
Total samples 24884 20798 19264 7252 4881 4776 4851 4627 4630 4312 3910
Avg. tokens 958 943 984 870 1656 1601 1623 919 901 917 880
Seed source Industry reports & expert reviewes Clinical databases Platform analytics
The LLMs referenced in Figure 7 are fine-tuned using QLoRA [19], configured with a LoRA rank of 8 and trained for 3 epochs, on two H100 GPUs, with models quantized to 4-bit precision via bitsandbytes. Other open-source models are fine-tuned for 10 epochs using LoRA [25]. The inference times reported in Figure 5 and Table 8 are benchmarked by deploying the open-source models on a single A40 GPU to mitigate network variability.
@@
PMC [68] Hierarchical Planning Low Middle Multiple
MedAgents [50] Discussion & Voting Low Middle Multiple
Debate(short) [20] Debate Low Middle Multiple
Debate(long) [20] Debate Low Middle Multiple
CoA [72] Chain Collaboration Low Middle Multiple
@@
The appendix provides a more extensive report on experimental results across additional metrics. These include BERTScore-P, BERTScore-R, BLEU-4 [41] for measuring text alignment, METEOR [6] which considers semantically equivalent phrases, ROUGE-2 for assessing content breadth, and ROUGE-L [35] which evaluates deep semantic alignment, particularly for long texts.
Appendix C More Experiments and Analysis
C.1 Efficiency and Cost Analysis
Table 8 presents a comparative analysis of Cochain against other multi-agent systems, focusing on key operational metrics: Inference Time, Token Throughput, and Cost. A noteworthy observation on the Automotive and E-commerce datasets is that Cochain’s average input token count is approximately one-twentieth of that utilized by PMC and CoA. This disparity suggests that these alternative systems expend a considerable volume of tokens on inter-agent communication. In contrast, Cochain innovatively transforms such communication into structured knowledge graphs and prompts tree, a mechanism that effectively curtails operational costs. The overall results compellingly demonstrate that Cochain achieves substantial reductions in inference time, token throughput, and aggregate cost when benchmarked against other multi-agent approaches. Remarkably, the efficiency metrics for Cochain closely approximate those observed for single-agent systems, and its average cost is even lower, underscoring a significant advantage of our proposed framework in terms of computational resource utilization and operational expenditure.
Table 8: Comparative analysis of inference time, token throughput, and cost across three datasets. Inference time is measured using a locally deployed Qwen2-7B model to obviate network latency associated with API calls. Conversely, token throughput and cost metrics are derived using Claude-3.5-Haiku.

### v2

@@ 4.2 Result
Figure 6: Stage-wise gains of Cochain over fine-tuning on the Automotive workflow ($\Delta=\text{+Cochain}-\text{Finetuned}$). (a) Jointly fine-tuned backbone (Qwen2-7B). (b) Stage-wise fine-tuned backbones. Bars show improvements in BS-F, GLEU, and ROUGE-L. The box reports the overall GPT-4 score gain of the integrated solution.
@@
Table 5: Result of ablation study. We highlight the largest and second-largest performance drops caused by removing each component.
4.3 Key Design Factors
@@
To validate Cochain’s effectiveness, we conduct ablation experiments with three variants: (1) w/o Knowledge Graph; (2) w/o Causal Chain; and (3) w/o Prompts Tree. As shown in Table 5, all variants exhibit performance degradation across all metrics, confirming each component’s importance. Notably, removing the prompts tree causes the most significant drop, highlighting its critical role in collaboration. The knowledge graph and causal chain components also prove essential, demonstrating their complementary functions in the reasoning process.
Hyperparameter Influence.
We vary two hyperparameters: the prompt count and the retrieval nodes top-$n$ from the knowledge graph. Figure 4 shows an inverted U-shape when fixing top-$n$ and increasing prompt count, indicating an optimal prompt budget rather than monotonic gains. When fixing the prompt budget and varying top-$n$, the trend is not monotonic. Overall, performance is more sensitive to prompt budgeting than simply retrieving more nodes, consistent with Table 5 where Prompts Tree yields the largest contribution.
Stage Dependency.
Table 4 shows that skipping any stage consistently degrades all metrics, indicating that each stage contributes non-redundant constraints for end-to-end workflow reasoning.
Figure 7: Cost comparison (Tokens, API Fees, Time) across baselines on pharmaceutical dataset. Cochain’s total costs include the one-time Prompts Tree setup and costs for varying evaluation queries.
4.4 Cross-Domain and Specialized Settings
We evaluate Cochain on the Automotive workflow under two regimes: a jointly fine-tuned backbone trained on the union of all stage data, and stage-wise fine-tuned backbones specialized per stage. Figure 6 reports stage-wise gains over the corresponding fine-tuned baselines, where $\Delta=\text{+Cochain}-\text{Finetuned}$. With joint fine-tuning (Figure 6(a)), Cochain improves all stages and metrics, with larger gains in later stages where downstream constraints are more tightly coupled with upstream decisions, indicating better constraint propagation without direct, token-intensive interaction. With stage-wise fine-tuning (Figure 6(b)), Cochain consistently boosts each specialized backbone, showing it can coordinate domain-expert agents into a coherent workflow solution. The integrated four-stage outputs also receive higher GPT-4 scores in both regimes, supporting end-to-end quality improvements beyond stage-local changes.
4.5 Efficiency and Cost Analysis
Figure 7 breaks down efficiency into one-time setup overhead and per-query running cost. In contrast to interaction-heavy baselines, Cochain shifts most computation to an offline setup stage and keeps inference lightweight. This yields consistent savings across budgets: for 1000 evaluation queries, Cochain consumes only 1.98M tokens in total, compared with 8.94M for CoA and 8.08M for Debate(long). The same trend holds in monetary cost, where Cochain stays at \$5.80 versus \$18.20 and \$20.12, and in wall-clock time, where Cochain finishes in 79,923s versus 253,184s and 574,949s. While the setup stage introduces an upfront cost, it is paid once and then reused, so the marginal cost grows slowly with more queries and the total advantage widens at scale. This shows that Cochain delivers high-quality solutions at low cost, effectively alleviating excessive collaboration.
@@ 5 Conclusion
In this paper, we propose Cochain, a collaboration prompting framework that balances Insufficient Collaboration and Excessive Collaboration in LLM agent workflows. Cochain bridges single-agent depth and multi-agent breadth via reusable collaboration artifacts, achieving low operational cost, fast inference, strong interpretability, and robust end-to-end reasoning. It builds a Prompts Tree to distill and propagate stage-wise constraints, enabling coherent multi-stage solutions without token-heavy interactions. Cochain also constructs a cross-stage Collaborative Knowledge Graph with tacit knowledge elicited through counterfactual querying, which strengthens relational reasoning and reduces hallucinations. In addition, the causal chain mechanism exposes explicit dependency traces, improving transparency and controllability. Extensive experiments on specialized and cross-domain settings show consistent improvements in scenarios affected by both collaboration extremes, and expert evaluation further indicates that smaller models equipped with Cochain can outperform GPT-4 on workflow tasks.
Appendix A Algorithms and Examples to Better Understand Cochain
Algorithm 1 Collaborative Knowledge Graph Construction
1:Datasets $D=\{D_{1},\dots,D_{N}\}$; LLM agents $A=\{A_{1},\dots,A_{N}\}$; Evaluation LLM $E$; Stage Labels $L=\{L_{1},\dots,L_{N}\}$.
2:A collaborative knowledge graph $\mathrm{KG}_{\text{collab}}$.
3:Initialize: $\mathrm{KG}_{\text{explicit}}\leftarrow\emptyset$, $\mathrm{KG}_{\text{tacit}}\leftarrow\emptyset$
4:
[…]
6:for $i\leftarrow 1$ to NN do
7:  for all $(\text{instruction},\text{response})\in D_{i}$ do
8:   $\mathrm{triples}\leftarrow\text{ExtractTriples}(\text{instruction},\text{response})$
9:   $\text{LabelNodesInTriples}(\mathrm{triples},L_{i})$
10:   $\mathrm{KG}_{\text{explicit}}\leftarrow\mathrm{KG}_{\text{explicit}}\cup\mathrm{triples}$
11:  end for
12:end for
13:
[…]
15:for $i\leftarrow 1$ to NN do
16:  for all $(\text{instruction},\text{response})\in D_{i}$ do
17:   $\mathrm{cf\_instruction}\leftarrow\text{GenerateCounterfactual}(\text{instruction})$
18:   repeat
19:   $\mathrm{cf\_response}\leftarrow A_{i}.\text{query}(\mathrm{cf\_instruction})$
20:   $(\text{evaluation},\text{feedback})\leftarrow E.\text{evaluate}(\mathrm{cf\_response})$
21:   if $\text{evaluation}\neq\text{"reasonable"}$ then
22:     $\mathrm{cf\_instruction}.\text{append}(\text{feedback})$
23:   end if
24:   until $\text{evaluation}=\text{"reasonable"}$
25:   $\mathrm{tacit\_triples}\leftarrow\text{ExtractTriples}(\mathrm{cf\_instruction},\mathrm{cf\_response})$
26:   $\text{LabelNodesInTriples}(\mathrm{tacit\_triples},L_{i})$
27:   $\mathrm{KG}_{\text{tacit}}\leftarrow\mathrm{KG}_{\text{tacit}}\cup\mathrm{tacit\_triples}$
28:  end for
29:end for
30:
[…]
33:return $\mathrm{KG}_{\text{collab}}$
Algorithm 2 Causal Chain Retrieval and Construction
1:User query $\mathrm{query}$; Stage-labeled knowledge graph $\mathrm{KG}_{\text{collab}}$; Text encoder $\mathrm{Encoder}$; Seed nodes count $\mathrm{top\_n}$; Chain length limit $\mathrm{max\_depth}$.
2:A textual causal chain $\mathrm{causal\_chain\_text}$.
3:
[…]
11:
[…]
14:for each $\mathrm{start\_node}$ in $\mathrm{seed\_nodes}$ do
15:  $\triangleright$ Search paths crossing stages via bridge nodes with multiple labels
16:  $\mathrm{paths}\leftarrow\mathrm{KG}_{\text{collab}}.\text{FindCrossStagePaths}(\mathrm{start\_node},\mathrm{min\_depth}\leftarrow 2,\mathrm{max\_depth})$
17:  $\mathrm{all\_paths}\leftarrow\mathrm{all\_paths}\cup\mathrm{paths}$
18:end for
19:
[…]
22:return $\mathrm{causal\_chain\_text}$
Algorithm 3 Prompts Tree Construction
1:Workflow stages $S=\{S_{1},\dots,S_{n}\}$; Domain agents $A=\{A_{1},\dots,A_{n}\}$; Seed Q-A pair $\mathrm{InitialSeedQA}$; Best prompts count $m$.
2:A prompt tree $\mathrm{PromptsTree}$.
3:Initialize:
[…]
7:while $\mathrm{queue}\neq\emptyset$ do
8:  $\mathrm{parentNode},\mathrm{currentAnswer},\mathrm{currentStage}\leftarrow\mathrm{queue}.\text{pop\_front}()$
9:  $\triangleright$ Phase 1: Distill Prompts
10:  $\mathrm{distilledPrompts}\leftarrow A_{\mathrm{currentStage}}.\text{DistillPrompts}(\mathrm{currentAnswer})$
11:  $\triangleright$ Phase 2: Select Best Prompts
12:  $\mathrm{bestPrompts}\leftarrow A_{\mathrm{currentStage}}.\text{SelfEvaluate}(\mathrm{distilledPrompts},m)$
13:  $\triangleright$ Phase 3: Grow Tree
14:  if $\mathrm{currentStage}\neq S_{n}$ then $\triangleright$ Proceed if not the last stage
15:   $\mathrm{nextStage}\leftarrow S_{\text{index}(\mathrm{currentStage})+1}$
16:   for each $\mathrm{promptText}$ in $\mathrm{bestPrompts}$ do
17:     $\mathrm{childNode}\leftarrow\text{CreateNode}(\mathrm{promptText})$
18:     $\mathrm{parentNode}.\text{add\_child}(\mathrm{childNode})$
19:     $\mathrm{newQuestion}\leftarrow\text{GenerateQuestionForNextStage}(\mathrm{promptText},\mathrm{nextStage})$
20:     $\mathrm{newAnswer}\leftarrow A_{\mathrm{nextStage}}.\text{query}(\mathrm{newQuestion})$
21:     $\mathrm{queue}.\text{push\_back}((\mathrm{childNode},\mathrm{newAnswer},\mathrm{nextStage}))$
22:   end for
23:  end if
24:end while
25:return $\mathrm{PromptsTree}$
As shown in the Figure 9, after applying Cochain, the agent anticipates that carbon fiber requires complex molding processes, completely different from traditional metal stamping, and that qualified suppliers are scarce, thus requiring consideration of supply chain stability and high procurement costs significantly impacting vehicle pricing and profit margins. The prompts tree anticipates existing production processes, guiding the agent to consider compatibility constraints between new equipment and existing stamping-welding lines. Based on this, Cochain generates superior answers compared to single agents.
Figure 8: Examples of implementing a core component in the Cochain framework.
Figure 9: Examples of single agent and complete workflow of Cochain
Appendix B Proof: Tacit Knowledge & Counterfactual Output Model
@@
$P(\tilde{Y}_{ij}\mid\tilde{X}_{ij})=\int_{\Theta}P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})P(h_{ij}\mid\theta_{ij})P(\theta_{ij}\mid\tilde{X}_{ij})d\theta_{ij}$ (10)
@@
$P(\tilde{Y}_{ij}\mid\tilde{X}_{ij})=\int_{\Theta}P(\tilde{Y}_{ij},\theta_{ij}\mid\tilde{X}_{ij})d\theta_{ij}$
@@
$P(\tilde{Y}_{ij}\mid\tilde{X}_{ij})=\int_{\Theta}P(\tilde{Y}_{ij}\mid\theta_{ij},\tilde{X}_{ij})P(\theta_{ij}\mid\tilde{X}_{ij})d\theta_{ij}$ (11)
@@
$P(\tilde{Y}_{ij}\mid\theta_{ij},\tilde{X}_{ij})=\int P(\tilde{Y}_{ij},h_{ij}\mid\theta_{ij},\tilde{X}_{ij})dh_{ij}$
$\int P(\tilde{Y}_{ij}\mid h_{ij},\theta_{ij},\tilde{X}_{ij})P(h_{ij}\mid\theta_{ij},\tilde{X}_{ij})dh_{ij}$
@@
$P(\tilde{Y}_{ij}\mid\tilde{X}_{ij})=\int_{\Theta}P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})P(h_{ij}\mid\theta_{ij})P(\theta_{ij}\mid\tilde{X}_{ij})d\theta_{ij}$
@@
Appendix C Implementation Details
Definition of Business Workflow: It is a complex process with many interconnected stages. These stages are specialized. The key feature is that these stages strongly depend on each other. A decision in one stage requires more than just expert knowledge of that area. It must also anticipate and use knowledge and constraints from the other stages.
@@ Appendix C Implementation Details
The LLMs referenced in Figure 5 (Main Paper) are fine-tuned using QLoRA Dettmers et al. [2023], configured with a LoRA rank of 8 and trained for 3 epochs, on two H100 GPUs, with models quantized to 4-bit precision via bitsandbytes. Other open-source models are fine-tuned for 10 epochs using LoRA Hu et al. [2022]. The inference times reported in Figure 10 and Table 7 are benchmarked by deploying the open-source models on a single A40 GPU to mitigate network variability.
@@
PMC Zhang et al. [2025a] Hierarchical Planning Low Middle Multiple
MedAgents Tang et al. [2024] Discussion & Voting Low Middle Multiple
Debate(short) Du et al. [2024] Debate Low Middle Multiple
Debate(long) Du et al. [2024] Debate Low Middle Multiple
CoA Zhang et al. [2024b] Chain Collaboration Low Middle Multiple
@@
The appendix provides a more extensive report on experimental results across additional metrics. These include BERTScore-P, BERTScore-R, BLEU-4 Papineni et al. [2002] for measuring text alignment, METEOR Banerjee and Lavie [2005] which considers semantically equivalent phrases, ROUGE-2 for assessing content breadth, and ROUGE-L Lin [2004] which evaluates deep semantic alignment, particularly for long texts.
Appendix D More Experiments and Analysis
D.1 More Efficiency and Cost Analysis
Table 7 presents a comparative analysis of Cochain against other multi-agent systems, focusing on key operational metrics: Inference Time, Token Throughput, and Cost. A noteworthy observation on the Automotive and E-commerce datasets is that Cochain’s average input token count is approximately one-twentieth of that utilized by PMC and CoA. This disparity suggests that these alternative systems expend a considerable volume of tokens on inter-agent communication. In contrast, Cochain innovatively transforms such communication into structured knowledge graphs and prompts tree, a mechanism that effectively curtails operational costs. The overall results compellingly demonstrate that Cochain achieves substantial reductions in inference time, token throughput, and aggregate cost when benchmarked against other multi-agent approaches. Remarkably, the efficiency metrics for Cochain closely approximate those observed for single-agent systems, and its average cost is even lower, underscoring a significant advantage of our proposed framework in terms of computational resource utilization and operational expenditure. Table 8 presents the retrieval contribution of each component during the inference process. The prompt tree retrieval employs a depth-first search (DFS) strategy. In terms of time complexity, this is equivalent to a simple path traversal with a complexity of $O(D)$, where $D$ denotes the depth of the business process. Figure 10 shows that as reasoning rounds increase, the time gap between thinking and non-thinking models grows dramatically for PMC but remains moderate for Cochain, demonstrating Cochain’s superior efficiency as complexity increases.
Figure 10: ExampReasoning time differential.
Table 7: Comparative analysis of inference time, token throughput, and cost. Inference time is measured using a locally deployed Qwen2-7B model to obviate network latency associated with API calls. Conversely, token throughput and cost metrics are derived using Claude-3.5-Haiku.